PLease show in detail how to solve the following problem We are getting stuck on the fraction portion. d/4 - 8 = -5 and t/5 + 6 = -2
Answer by checkley75(3666)

D/4-8=-5
D/4=-5+8
D/4=3 NOW WE HAVE TO CROSS MULTIPLY D=4*3
D=12 ANSWER.
PROOF
12/4-8=-5
3-8=-5
-5=-5
---------------------------------------------------------
T/5+6=-2
T/5=-2-6
T/5=-8 AGAIN CROSS MULTIPLY
T=5*-8
T=-40 ANSWER.
PROOF
-40/5+6=-2
-8+6=-2
-2=-2